Can you help me...Solve this equation by using the square root property?

(x-1)^2=81.....FYI ^2 means "square"

Answer by checkley71(8403)

(x-1)^2=81
(x-1)^2=9^2 take the sqrt of each side.
x-1=9
x=10 answer.
proof
(10-1)^2=81
9^2=81
81=81